40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1185     Entry time: Mon Dec 8 00:10:42 2008
Author: caryn 
Type: Summary 
Category: General 
Subject: calibrating the jenne laser 
I apologize in advance for the long list of numbers in the attachment. I can't seem to make them hide for some reason.

So, since Jenne's laser will probably be used for the Stoch mon calibration, Alberto and I took some measurements to calibrate Jenne's laser.
We focused the beam onto the New Focus RF 1GHz photodetector that we stole from rana's lab (powered with NewFocus power 0901). Measured the DC output of the photodetector with scope. Aligned the beam so DC went up (also tried modulating laser at 33MHz and aligning so 33MHz peak went up). Hooked up the 4395a Spectrum/Network Analyzer to the laser and to the AC out of the photodetector (after calibrating Network analyzer with the cables) so that the frequency response of the laser*photodetector could be measured.
(Note: for a while, we were using a splitter, but for the measurements here, I got rid of the splitter and just sent the RFout through the cables to channel A for the calibration, sent RFout to the laser and photodetector to channel A for the measurement)

Measured the frequency response. At first, we got this weird thing with a dip around 290MHz (see jcal_dip_2_norm.png below).
After much fiddling, it appeared that the dip was from the laser itself. And if you pull up just right on the corner of this little metal flap on the laser (see picture), then the dip in the frequency response seems to go away and the frequency response is pretty flat(see jcal_flat_3_norm below). If you press down on the flap, the dip returns. This at least happened a couple of times.
Note that despite dividing the magnitude by the DC, the frequency responses don't all line up. I'm not sure why. In some cases the DC was drifting a bit(I presume the laser was coming out of alignment or decided to align itself better) and maybe with avgfactor=16, and measuring mean DC on the scope, it made the DC meas not match up the the frequ resp meas...
I've attached the data for the measurements made (I'm so sorry for all the #'s. I can't figure out how to hide them)
name/lasercurrent/DC/analyzer SourcePower/analyzer avgfactor
jcal7_1/I=31.7mA/DC=-4.41/SourcePower=0dBm/avgfactor=16
jcal7_2/I=31.7mA/DC=-1.56/SourcePower=0dBm/avgfactor=none
jcal8_1/I=31.7mA/DC=-4.58/SourcePower=0dBm/avgfactor=16
jcal8_2/I=31.7mA/DC=-2.02/SourcePower=0dBm/avgfactor=16
jcal8_3/I=31.7mA/DC=-3.37/SourcePower=0dBm/avgfactor=16
Note also that the data from the 4395a seems to have column1-frequency, column2-real part, column3-imaginary part...I think. So, to calculate the magnitude, I just took (column2)^2+(column3)^2.


To get sort of an upper-bound on the DC, I measured how DCmax varied with laser current, where DCmax is the DC for the best alignment I could get. After setting the current, the laser was modulated at 33MHz and the beam was aligned such that the 33MHz peak in the photodetector output was as tall as I could manage. Then DC was measured. See IvsDCmax.png. Note the DC is negative. I don't know why.

Also, the TV's don't look normal, the alarm's going off and I don't think the mode cleaner's locked.
Attachment 1: IvsDCmax.png  16 kB  | Hide | Hide all
IvsDCmax.png
Attachment 2: data.tar.gz  15 kB  Uploaded Mon Dec 8 10:17:14 2008
Attachment 3: jcal_dip_2_norm_log.png  14 kB  Uploaded Tue Dec 9 11:00:35 2008  | Hide | Hide all
jcal_dip_2_norm_log.png
Attachment 4: jcal_flat_3_norm_log.png  15 kB  Uploaded Tue Dec 9 11:00:59 2008  | Hide | Hide all
jcal_flat_3_norm_log.png
ELOG V3.1.3-